OpenStudy (anonymous):

What is the solution for: 2x-5y=-1 3x-4y=2

OpenStudy (anonymous):

Simplifying (13x + -4y) + (2x + 5y) = 0 Remove parenthesis around (13x + -4y) 13x + -4y + (2x + 5y) = 0 Remove parenthesis around (2x + 5y) 13x + -4y + 2x + 5y = 0 Reorder the terms: 13x + 2x + -4y + 5y = 0 Combine like terms: 13x + 2x = 15x 15x + -4y + 5y = 0 Combine like terms: -4y + 5y = 1y 15x + 1y = 0 Solving 15x + 1y = 0 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-1y' to each side of the equation. 15x + 1y + -1y = 0 + -1y Combine like terms: 1y + -1y = 0 15x + 0 = 0 + -1y 15x = 0 + -1y Remove the zero: 15x = -1y Divide each side by '15'. x = -0.06666666667y Simplifying x = -0.06666666667y